Which expression is equal to real number? a) (-4i)^11. b) (4+5i) (4-5i) c) (2+3i)^2. d) (6+8i) (8+6i)

Grade:9

1 Answers

Arun
6 years ago
(-4 i)11 = 411 i11 = -411 i = imaginary
(4+5i)(4-5i) = 42 - 52 i2 = 16+25 = 41 = real number
(2+3i)2 = -5 + 12i = complex number with imaginary part
(6+8i)(8+6i)= 48 - 48 + 100i = 100 i = imaginary
 
hence optuon 'b' is equal to real number.